A methods and measurement analyst wants to develop a time standard for a certain task. In a preliminary study, he observed one worker perform the task six times with an average observed time of 25 seconds and a standard deviation of 3.8 seconds.

What is the standard time for this task if the employee worked at a 15.2 percent faster pace than average, and an allowance of 20.6 percent of the workday is used?

(Keep one decimal place in your answer)
Answer #1

Average Time = 25 Seconds

Performance Rating = 1+0.152 = 1.152

Normal Tine = Average Time*Performance Rating

Normal Time = 25*1.152 = 28.8 Seconds

Standard Time = Normal Time*(1+Allowance Factor)

Standard Time = 28.8*(1+0.206)

Standard Time = 34.7328 seconds

Standard Time = 34.7 seconds